Compute the mean and standard deviation of the population 2,6,8,10,12,14


1
Expert's answer
2021-04-27T17:07:21-0400

"X = \{2,6,8,10,12,14\}"

n=7 because there are 7 number in the population.

The mean is

"\displaystyle \mu = E[X] = \frac{\sum X_i}{n} = \frac{2+6+8+10+12+14}{7}=7.429"

The standard deviation for population is

"\displaystyle s = \sqrt{\frac{\sum (X_i - \mu)^2}{n}}= 3.944"


Answer: 7.429, 3.944